OpenStudy (anonymous):

solve each equation on the interval [0,2pi) 2 sin theta cos theta = -1

OpenStudy (kc_kennylau):

"each"?

OpenStudy (kc_kennylau):

do you remember the sine of double angle theorem? :)

OpenStudy (anonymous):

*the equation

OpenStudy (anonymous):

sin2x = 2sinxcosx

OpenStudy (kc_kennylau):

the theorem is \(\sin(2\theta)=2\sin\theta\cos\theta\)

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

so you can solve it now? :)

OpenStudy (anonymous):

but how do i isolate theta?

OpenStudy (anonymous):

sin2x = -1

OpenStudy (kc_kennylau):

what are the values of 2x?

OpenStudy (anonymous):

can i just divide what do you mean?

OpenStudy (anonymous):

*divide out the 2

OpenStudy (kc_kennylau):

I mean what are the possible values of 2x if sin(2x)=-1? :)

OpenStudy (anonymous):

so x is 5pi/6 or pi/6?

OpenStudy (kc_kennylau):

nope :/

OpenStudy (anonymous):

so how do i get the value of x then?

OpenStudy (kc_kennylau):

find the value of 2x first. What will be the range of 2x? :)

OpenStudy (anonymous):

-1

OpenStudy (kc_kennylau):

the range of x is \([0,2\pi)\), so the range of 2x is?

OpenStudy (anonymous):

[0,2pi)?

OpenStudy (kc_kennylau):

nvm just find the values of 2x if sin(2x)=-1

OpenStudy (kc_kennylau):

sine of what is negative one?

OpenStudy (anonymous):

-90

OpenStudy (kc_kennylau):

and a few more?

OpenStudy (kc_kennylau):

(remember it should be in the interval \([0,4\pi)\))

OpenStudy (anonymous):

-pi/2?

OpenStudy (kc_kennylau):

\([0,4\pi)\) means that it's greater than 0 and smaller than 4pi

OpenStudy (anonymous):

3pi/2

OpenStudy (kc_kennylau):

and more?

OpenStudy (anonymous):

isn't that it for -1?

OpenStudy (kc_kennylau):

what about \(\frac72\pi\)

OpenStudy (anonymous):

oh i don't have that on my rad circle

OpenStudy (kc_kennylau):

your circle is only to \(2\pi\). Since x is \([0,2\pi)\), 2x is \([0,4\pi)\).

OpenStudy (kc_kennylau):

so now \(2x\) is \(\frac{3\pi}2\) or \(\frac{7\pi}2\)

OpenStudy (kc_kennylau):

what would \(x\) be? :)

OpenStudy (anonymous):

3pi/4 or 7pi/4

OpenStudy (kc_kennylau):

well done :)

OpenStudy (anonymous):

thank you soo much!!

OpenStudy (kc_kennylau):

no problem :)